# Einstein connection of nonsymmetric Riemannian manifold

### 04 Mar 2026 v2

@@ Abstract
A. Einstein considered a nonsymmetric (0,2)-tensor $G=g+F$, where $g$ is a pseudo-Riemannian metric and $F\neq 0$ is skew-symmetric, and a linear connection $\nabla$ with torsion $T$ such that $(\nabla_{X}\,G)(Y,Z)=-G(T(X,Y),Z)$. M. Prvanović (1995) obtained the explicit form of the Einstein connection of an almost Hermitian manifold. In this paper, first, we present the result above in coordinate-free form, and then extend it to almost contact metric manifolds satisfying the $f^{2}$-torsion condition $T(f^{2}X,Y)=T(X,f^{2}Y)=f^{2}T(X,Y)$, where $F(X,Y)=g(X,fY)$. We then derive the Einstein connection of nonsymmetric pseudo-Riemannian, in particular, weak almost Hermitian manifolds $(M,f,g)$, satisfying the $f^{2}$-torsion condition, give explicit formulas for the torsion in terms of $\nabla^{g}F$, $dF$ and a new (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$, and show that in the almost Hermitian case, our results reduce to the coordinate-free form of Prvanović’s solution. Finally, we discuss special Einstein connections, and indicate the Gray-Hervella classes. Illustrative examples are given, including the construction of a weighted product.
Keywords: nonsymmetric pseudo-Riemannian manifold, weak almost Hermitian manifold, almost contact metric structure, Einstein connection, $f^{2}$-torsion condition.
@@ 1 Introduction
A. Einstein worked on various versions of the Unified Field Theory (Nonsymmetric Gravitational Theory – NGT), see [2]. This theory was intended to unite the gravitation theory, to which General Relativity is related, and the theory of electromagnetism. Beginning in 1950, A. Einstein used a real nonsymmetric basic (0,2)-tensor $G$, whose symmetric part $g$ is associated with gravity, and the skew-symmetric part $F$ is associated with electromagnetism. In his attempt to construct a Unified Field Theory, (briefly NGT), A. Einstein [2] considered a snooth manifold $(M,G=g+F)$ with a linear connection $\nabla$ with torsion $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$ satisfying the following condition:
@@
called an Einstein connection. The idea of a nonsymmetric basic tensor was revisited by J.W. Moffat [7], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [6] point to potential improvements and challenges within NGT. M.I. Wanas et al. [11] studied path equations of a particle in Einstein’s nonsymmetric geometry. Recent approaches to modified gravity often rely on differential geometry, including torsion and non-metricity, as a natural extension of General Relativity. Connections with totally skew-symmetric torsion, i.e., $T(X,Y,Z)$ is a 3-form, satisfying T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), are important due to the relations with mathematical physics, as they admit a coupling to spinor fields and lead to holonomy and rigidity phenomena. In this case, the dimension of the space of parallel spinors coincides with the number of preserved supersymmetries. Such connections arise naturally in supersymmetric string theories and non-linear $\sigma$-models, as well as in gravitational models [4]. S. Ivanov and M. Lj. Zlatanović (see [5]) presented conditions for the existence and uniqueness of the Einstein connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression using an almost contact metric structure. In this case, the torsion is determined by $dF$: $T(X,Y,Z)=-\frac{1}{3}\,dF(X,Y,Z)$, and the difference tensor is given by
@@
A linear connection $\nabla$} on a Riemannian manifold $(M,g)$ is said to be statistical if it is torsionless and tensor $({\nabla}g)(X,Y,Z):=({\nabla}_{X}\,g)(Y,Z)$ is symmetric in all its entries, equivalently, the cubic form $K(X,Y,Z):=g(K(X,Y),Z)$ (associated with the difference tensor (2.9) of $\nabla$) is symmetric. The statistical structure has applications in theory of probability and statistics as well as in information geometry. Unfortunately, we cannot directly apply the statistical structure to Einstein’s nonsymmetric geometry: if the cubic form $K(X,Y,Z)$ is symmetric, then by (2.11) and (2.12), $T=K=0$ is true; hence, $\nabla=\nabla^{g}$.
Weak metric structures, developed by V. Rovenski et. al, see [9], generalize the classical metric $f$-structures and are well suited for a study of $G=g+F$ in NGT with an arbitrary skew-symmetric tensor $F$. A weak almost Hermitian manifold $(M,f,g)$ is a Riemannian manifold $(M,g)$ of dimension $2m\geq 4$ endowed with non-singular skew-symmetric endomorphism $f$ and the 2-form $F$ such that $F(X,Y)=g(X,fY)$, see [12]. V. Rovenski and M. Zlatanović [12, 10] were the first to apply the weak $f$-structure, in particular, the weak almost contact structure, to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$.
M. Prvanović [8] obtained the explicit form of the Einstein connection of almost Hermitian manifold, and using its torsion tensor, characterized the sixteen classes (by A. Gray and L.M. Hervella [3]) of almost Hermitian manifold.
In this paper, we obtain the explicit form of the Einstein connection of an NGT space, in particular, a weak almost Hermitian manifold, satisfying some natural condition, called the $f^{2}$-torsion condition. This condition disappears for the classical case, hence our result generalizes the result by M. Prvanović. We also present the explicit form of the Einstein connection of an almost contact metric manifold satisfying the $f^{2}$-torsion condition.
The paper has an Introduction and six sections. In Section 2 we review basics of NGT, give examples, prove several lemmas, and discuss the relation between torsion and the contorsion, $K$, of an Einstein connection emphasizing the role of $f^{2}$-torsion condition. In Section 3 we deduce equations for Einstein connection of an almost contact metric manifold with the $f^{2}$-torsion condition (Theorem 3.3). In Section 4 we generalize the result of M. Prvanović and deduce equations for an Einstein connection with the $f^{2}$-torsion condition of nonsymmetric pseudo-Riemannian manifold (Theorem 4.3), in particular, for weak almost Hermitian manifold (Corollary 4.5). Long expressions of Section 4 are given in Section 5.
@@ 2 Einstein’s nonsymmetric geometry
The metricity condition (1.1) can be writen in the form
@@
Example 2.3.
Let $(M,f,g)$ be a weak nearly Kähler manifold, considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$ with $F(X,Y)=g(X,fY)$. If $\nabla$ is an Einstein connection on $M$ with totally skew-symmetric torsion $T$, then the $f$-torsion condition holds:
$T(fX,Y)=T(X,fY)=-f\,T(X,Y),$ (2.17)
@@
The following skew-symmetry of the difference tensor $K$ is obtained using (2.12):
$K_{X}Y=-K_{Y}X,\quad\text{or, equivalently,}\quad K_{X}Y=\tfrac{1}{2}\,T(X,Y),$ (2.20)
and characterizes special Einstein connections, see [8]. This condition expresses the skew–symmetry of $K$ only in the first two arguments, and doesn’t imply the total skew–symmetry of $K(X,Y,Z)$, which requires, in addition, condition (2.18). In this case, torsion tensor becomes totally skew–symmetric and satisfies the $f$-torsion condition, see Lemma 2.5.
@@
We represent the solution (an explicit form of the Einstein connection of an almost Hermitian manifold) of M. Prvanović [8] in the coordinate-free form.
Theorem 2.8 (see Theorem 1 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a nonsymmetric Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$ Then an Einstein connection $\nabla$ on $M$ is given by
$-(\nabla^{g}_{Y}F)(Z,X)-(\nabla^{g}_{Z}F)(X,Y).$ (2.26)
In view of (2.8), the Einstein connection $\nabla$ is special, see definition (2.20), if and only if the following identity holds:
Corollary 2.9 (see Theorem 2 of [8]).
Let $(M^{2n},J,g)$ be an almost Hermitian manifold considered as a generalized Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,JY)$. Then the torsion of a special Einstein connection $\nabla$ on $M$, see (2.20), is given by
@@ 3 Einstein connection of an almost contact metric manifold
The next consequences of Theorem 3.3 explicitly present the special Einstein connection with the totally skew-symmetric torsion satisfying the $f^{2}$-torsion condition.
@@
$2\,T(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)-(\nabla^{g}_{fZ}F)(fX,Y)-(\nabla^{g}_{fY}F)(fX,Z).$ (3.24)
@@
By Theorem 3.3, we have $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$. The rest of the proof is the same as the proof of Corollary 2.9. ∎
Corollary 3.8.
Under the conditions of Theorem 3.3, let $\nabla$ be a special Einstein connection. Then the torsion $T$ is totally skew-symmetric if and only if (3.7) is true; in this case,
$T(X,Y,Z)=\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z)=\tfrac{1}{6}\,dF(X,Y,X).$ (3.25)
Proof.
Assume that the torsion $T$ of $\nabla$ is totally skew-symmetric. Applying T(X,Y,Z)=T(X,Z,Y)T(X,Y,Z)=-T(X,Z,Y) to (3.24), we obtain
@@
Substituting the identity (3.26) into (3.24) yields the first equality of (3.25). Taking $Z\to fZ$ and $Y\to fY$ in (3.26) and using the equalities (3.7), $(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z)$, see (3.4), and $\nabla^{g}_{\xi}\,F=\nabla^{g}_{X}F(\xi,\,\cdot)=0$, in view of (2.5), we get the second equality of (3.25):
$(\nabla^{g}_{X}F)(Y,Z)=(\nabla^{g}_{Y}F)(Z,X)=(\nabla^{g}_{Z}F)(X,Y)=\tfrac{1}{3}\,dF(X,Y,Z).$
Conversely, if (3.25) holds, then $T$ is totally skew-symmetric, since $(\nabla^{g}_{X}F)(Y,Z)$ is skew-symmetric in $Y,Z$. The conditions $\nabla^{g}_{\xi}\,F=0$ and $\nabla^{g}\,\xi=0$ follow from Theorem 3.3. ∎
@@ 4 Einstein connection of an NGT space
In this section, we present explicitly an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, in particular, of a weak almost Hermitian manifold.
A pseudo-Riemannian manifold $(M^{2n},g)$ endowed with a skew-symmetric (1,1)-tensor $f$ of maximal rank $2n$ ($f^{2}$ is not necessarily equal to I-{I}) is called a weak almost Hermitian manifold. If $\nabla^{g}J=0$, then such $(M^{2n},J,g)$ is called a weak Kähler manifold, see [9].
The $f^{2}$-torsion condition (3.3) in terms of a self-adjoint (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$ reads
$T(\widetilde{Q}X,Y,Z)=T(X,\widetilde{Q}Y,Z)=T(X,Y,\widetilde{Q}Z).$
@@
The following theorem generalizes Theorem 2.8, it is valid for arbitrary $F$ satisfying (3.3), and for a non-degenerate tensor $P={I}-f^{2}$ we can use it to completely determine $T$.
@@
Let $\nabla$ be an Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$, satisfying the $f^{2}$-torsion condition (3.3), where $g(X,fY)=F(X,Y)$, and let the tensor $P:={I}-f^{2}$ have rank $2n$. If $fX\neq 0$, then the torsion $T$ of $\nabla$ is given by
$2\,T(Y,Z,\,f^{6}X)=-(\nabla^{g}_{X}F)((f+f^{3})Y,fZ)-(\nabla^{g}_{Y}F)(f^{2}Z,X)-(\nabla^{g}_{Z}F)(f^{2}X,Y)$
$+(\nabla^{g}_{fX}F)(f^{3}Y,Z)+(\nabla^{g}_{fX}F)(f^{2}Y,fZ)-(\nabla^{g}_{fY}F)(f^{3}Z,X)-(\nabla^{g}_{fZ}F)(f^{2}X,fY)$
$-(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,fY,Z)+(\nabla^{g}F)(P^{-1}(3\,\widetilde{Q}+2\,\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$-(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,fY,Z)+(\nabla^{g}F)(P^{-1}\widetilde{Q}f^{4}X,Y,fZ)+(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)$
$-\tfrac{1}{2}\,dF(Y,Z,(4\,\widetilde{Q}+3\,\widetilde{Q}^{2})f^{2}X)+\tfrac{1}{2}\,dF(Y,fZ,\widetilde{Q}f^{3}X)+\tfrac{1}{2}\,dF(fY,Z,\widetilde{Q}f^{3}X)$
$+dF(fY,fZ,(\widetilde{Q}+\widetilde{Q}^{2})X),$ (4.4)
where $\widetilde{Q}=-f^{2}-{I}$, and if $fX=0$, then
$T(Y,Z,X)=2(\nabla^{g}_{X}\,F)(Y,Z)+dF(Y,Z,X).$ (4.5)
@@
For $fX\neq 0$, it is carried out in a similar, but more computationally complex way than the proof of Theorem 2.8. Applying $Y\to fY$ and $Z\to fZ$ to (2.7), we have

### 24 Mar 2026 v3

# Einstein connection of nonsymmetric pseudo-Riemannian manifold

@@ Abstract
A. Einstein considered a linear connection $\nabla$ with torsion $T$ on a smooth manifold equipped with a nonsymmetric (0,2)-tensor $G=g+F$, where $g$ is a pseudo-Riemannian metric associated with gravity, and $F\neq 0$ is a skew-symmetric tensor associated with electromagnetism, such that $(\nabla_{X}\,G)(Y,Z)=-G(T(X,Y),Z)$. In this paper, we explicitly present the Einstein connection of a nonsymmetric pseudo-Riemannian manifold with non-degenerate $F$, satisfying the $f^{2}$-torsion condition $T(f^{2}X,Y)=T(X,f^{2}Y)=f^{2}T(X,Y)$, where $g(X,fY)=F(X,Y)$, and show that in the almost Hermitian case, it reduces to the M. Prvanović’s (1995) solution. We also discuss special Einstein connections, explicitly present the Einstein connection of almost contact metric manifolds satisfying the $f^{2}$-torsion condition, and give example in terms of weighted product of almost Hermitian manifolds.
Keywords: nonsymmetric pseudo-Riemannian manifold, Einstein connection, almost contact metric structure, $f^{2}$-torsion condition.
@@ 1 Introduction
A. Einstein worked on various versions of the Unified Field Theory (Nonsymmetric Gravitational Theory – NGT), see [2]. This theory was intended to unite the gravitation theory, to which General Relativity is related, and the theory of electromagnetism. Beginning in 1950, A. Einstein used a real nonsymmetric basic (0,2)-tensor $G$, whose symmetric part $g$ is associated with gravity, and the skew-symmetric part $F$ is associated with electromagnetism. In his attempt to construct a Unified Field Theory, (briefly NGT), A. Einstein [2] considered a smooth manifold $(M,G=g+F)$ with a linear connection $\nabla$ with torsion $T(X,Y)=\nabla_{X}Y-\nabla_{Y}X-[X,Y]$ satisfying the following condition:
@@
called an Einstein connection. The idea of a nonsymmetric basic tensor was revisited by J.W. Moffat [7], allowing for richer gravitational dynamics. Results by T. Jansen and T. Prokopec [6] point to potential improvements and challenges within NGT. M.I. Wanas et al. [11] studied path equations of a particle in Einstein’s nonsymmetric geometry. Recent approaches to modified gravity often rely on differential geometry, including torsion and non-metricity, as a natural extension of General Relativity. Connections with totally skew-symmetric torsion, i.e., $T(X,Y,Z)$ is a 3-form, satisfying T(X,Y,Z)=T(X,Z,Y)=T(Y,X,Z)T(X,Y,Z)=-T(X,Z,Y)=-T(Y,X,Z), are important due to the relations with mathematical physics, as they admit a coupling to spinor fields and lead to holonomy and rigidity phenomena. In this case, the dimension of the space of parallel spinors coincides with the number of preserved supersymmetries. Such connections arise naturally in supersymmetric string theories and non-linear $\sigma$-models, as well as in gravitational models [4]. S. Ivanov and M. Lj. Zlatanović (see [5]) presented conditions for the existence and uniqueness of the Einstein connection with totally skew-symmetric torsion on a manifold $(M,G=g+F)$ and gave its explicit expression using an almost contact metric structure. In this case, the torsion is determined by $T(X,Y,Z)=-\frac{1}{3}\,dF(X,Y,Z)$, and the difference tensor $K:=\nabla-\nabla^{g}$ is given by
@@
An important class of linear connections, having applications in theory of probability and statistics as well as in information geometry, are statistical connections, which are torsion-free and the tensor $({\nabla}g)(X,Y,Z):=({\nabla}_{X}\,g)(Y,Z)$ is symmetric with respect to all entries, equivalently, the cubic form $K(X,Y,Z):=g(K(X,Y),Z)$ (associated with the difference tensor (2.9) of $\nabla$) is symmetric. Unfortunately, we cannot directly apply the statistical structure to Einstein’s nonsymmetric geometry: if the cubic form $K(X,Y,Z)$ is symmetric, then by (2.11) and (2.12), $T=K=0$ is true; hence, $\nabla=\nabla^{g}$.
Weak metric structures, see [9], generalize the almost Hermitian and almost contact metric structures, as well as K. Yano’s $f$-structure, and are well suited for studying $(M,G=g+F)$ in NGT with an arbitrary skew-symmetric tensor $F$. A weak almost Hermitian manifold $(M,f,g)$ is a Riemannian manifold $(M,g)$ of even dimension $\geq 4$ endowed with non-singular skew-symmetric (1,1)-tensor $f$ and the 2-form $F(X,Y)=g(X,fY)$, see [12]. V. Rovenski and M. Zlatanović [12, 10] were the first to apply the weak almost Hermitian and weak almost contact structures to NGT of totally skew-symmetric torsion with $F(X,Y)=g(X,fY)$. M. Prvanović [8] obtained the explicit form of the Einstein connection of almost Hermitian manifold, and using its torsion tensor, characterized the sixteen classes (by A. Gray and L.M. Hervella [3]) of almost Hermitian manifold.
In this paper, we explicitly present the Einstein connection of an NGT space with non-degenerate $F$, that is, a weak almost Hermitian manifold, satisfying some natural condition, called the $f^{2}$-torsion condition. This condition disappears for the classical case, hence our result generalizes the result by M. Prvanović. We also explicitly present the Einstein connection (and the special Einstein connection defined by the property $K_{X}Y=-K_{Y}X$) of an almost contact metric manifold satisfying the $f^{2}$-torsion condition.
The paper has an Introduction and five sections. In Section 2 we review basics of NGT, give examples, prove several auxiliary statements, and discuss the relation between the torsion and contorsion of an Einstein connection emphasizing the role of $f^{2}$-torsion condition. In Section 3 we find the Einstein connection of an almost contact metric manifold with the $f^{2}$-torsion condition (Theorems 3.3 and 3.7). In Section 4 we present the Einstein connection of nonsymmetric pseudo-Riemannian manifold with non-degenerate $F$ and the $f^{2}$-torsion condition (Theorem 4.3), in particular, for weak almost Hermitian manifold (Corollary 4.5). Long expressions (terms $\delta_{i}$) of Section 4 are given in Section 5.
@@ 2 Einstein’s nonsymmetric geometry
The metricity condition (1.1) can be written in the form
@@
The following property $K_{X}Y=-K_{Y}X$ of the difference tensor $K$ (i.e., the skew–symmetry of $K(X,Y,Z)$ only in the first two arguments) by (2.11) reduces to
$T(Y,Z,fX)+T(X,Z,fY)=0,$ (2.19)
and characterizes special Einstein connections, i.e., the symmetric part of an Einstein connection $\nabla$ coincides with the Levi-Civita connection $\nabla^{g}$, see [8]. In this case, (2.12) yields $K_{X}Y=\tfrac{1}{2}\,T(X,Y)$. The condition (2.19) doesn’t imply the total skew–symmetry of $K(X,Y,Z)$, which, in addition, requires the condition (2.17).
@@ 3 Einstein connection of an almost contact metric manifold
The next results explicitly present the special Einstein connection of $(M^{2n+1},f,\xi,\eta,g)$ satisfying the $f^{2}$-torsion condition.
Theorem 3.7.
$2\,T(X,Y,Z)=-(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{fY}F)(fX,Z)+(\nabla^{g}_{fZ}F)(fX,Y).$ (3.24)
Proof.
Let $X,Y,Z$ be horizontal vector fields ($X,Y,Z\in\ker\eta$). Theorem 3.3 gives
$2(\nabla^{g}_{X}F)(Y,Z)=T(Y,fZ,fX)+T(fY,Z,fX).$ (3.25)
Using (2.19) (since $\nabla$ is a special Einstein connection), and $f^{2}=-I$ on $\ker\eta$, we obtain
$T(Y,fZ,fX)$ $=-T(X,fZ,fY)=T(fZ,X,fY)=-T(X,Y,Z),$
$T(fY,Z,fX)$ $=-T(X,Z,f^{2}Y)=T(X,Z,Y).$
Hence (3.25) becomes
$2(\nabla^{g}_{X}F)(Y,Z)=-T(X,Y,Z)+T(X,Z,Y).$ (3.26)
Next, replacing $X$ by $fZ$, $Y$ by $fX$, and $Z$ by $Y$ in (3.25), we get
$2(\nabla^{g}_{fZ}F)(fX,Y)$ $=T(fX,fY,f^{2}Z)+T(f^{2}X,Y,f^{2}Z)$
$=-T(fX,fY,Z)+T(X,Y,Z).$ (3.27)
Similarly, replacing $X$ by $fY$, $Y$ by $fX$, and $Z$ by $Z$ in (3.25), we obtain
$2(\nabla^{g}_{fY}F)(fX,Z)=-T(fX,fZ,Y)+T(X,Z,Y).$ (3.28)
Also, applying (2.19), we get
T(fX,fY,Z)+T(fX,fZ,Y)=T(fY,fX,Z)T(fZ,fX,Y)=0.T(fX,fY,Z)+T(fX,fZ,Y)=-T(fY,fX,Z)-T(fZ,fX,Y)=0. (3.29)
Subtracting (3) and (3.28) from (3.26), and using (3.29), we obtain
$2(\nabla^{g}_{X}F)(Y,Z)-2(\nabla^{g}_{fZ}F)(fX,Y)-2(\nabla^{g}_{fY}F)(fX,Z)$
$=\bigl(-T(X,Y,Z)+T(X,Z,Y)\bigr)-\bigl(-T(fX,fY,Z)+T(X,Y,Z)\bigr)$
$\qquad-\bigl(-T(fX,fZ,Y)+T(X,Z,Y)\bigr)$
$=-2\,T(X,Y,Z)+T(fX,fY,Z)+T(fX,fZ,Y)=-2\,T(X,Y,Z).$
Thus (3.24) is true for all horizontal vector fields $X,Y,Z$. If one of the vector fields is $\xi$, both sides vanish, hence the identity (3.24) holds for arbitrary vector fields on $M$. ∎
@@
Under the conditions of Theorem 3.7, the torsion $T$ (of a special Einstein connection $\nabla$) is totally skew-symmetric if and only if (3.7) is true; in this case,
$T(X,Y,Z)={\color[rgb]{1,0,0}-}\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z)={\color[rgb]{1,0,0}-}\tfrac{1}{6}\,dF(X,Y,Z).$ (3.30)
@@
Assume that the torsion $T$ is totally skew-symmetric. Then, by Theorem 3.7, we have (3.7) and (3.24). Since $T$ is totally skew-symmetric, using (3.24), we get
@@
Using this in (3.24), we obtain the first equality in (3.30): $T(X,Y,Z)=-\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z)$.
Replacing $Y$ by fYfY and $Z$ by $fZ$ in (3.31), and using $f^{2}=-I$ on $\ker\eta$, together with
$(\nabla^{g}_{X}F)(fY,fZ)=-(\nabla^{g}_{X}F)(Y,Z),$
we obtain
$(\nabla^{g}_{X}F)(Y,Z)=(\nabla^{g}_{Y}F)(Z,X)=(\nabla^{g}_{Z}F)(X,Y).$
Therefore,
$dF(X,Y,Z)=(\nabla^{g}_{X}F)(Y,Z)+(\nabla^{g}_{Y}F)(Z,X)+(\nabla^{g}_{Z}F)(X,Y)=3(\nabla^{g}_{X}F)(Y,Z);$
hence
$T(X,Y,Z)=-\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z)=-\tfrac{1}{6}\,dF(X,Y,Z).$
that completes the proof of (3.30). Conversely, assume that
$T(X,Y,Z)=-\tfrac{1}{2}(\nabla^{g}_{X}F)(Y,Z).$
Since $(\nabla^{g}_{X}F)(Y,Z)$ is skew-symmetric in $Y$ and $Z$, it follows that $T$ is skew-symmetric in $Y$ and $Z$; therefore, $T$ is totally skew-symmetric. ∎
@@ 4 Einstein connection of an NGT space
In this section, we explicitly present the Einstein connection of a nonsymmetric pseudo-Riemannian space $(M,G=g+F)$ with non-degenerate $F$, in particular, of a weak almost Hermitian manifold. Recall that a pseudo-Riemannian manifold $(M^{2n},g)$ endowed with a skew-symmetric (1,1)-tensor $f$ of maximal rank $2n$ ($f^{2}$ is not necessarily equal to I-{I}) is called a weak almost Hermitian manifold [9]. In this case, we introduce a new (1,1)-tensor $\widetilde{Q}=-f^{2}-{I}$, and the $f^{2}$-torsion condition (3.3) in terms of $\widetilde{Q}$ reads
$T(\widetilde{Q}X,Y,Z)=T(X,\widetilde{Q}Y,Z)=T(X,Y,\widetilde{Q}Z).$ (4.1)
@@
The following theorem generalizes [8, Theorem 1], and for a non-degenerate tensor $P={I}-f^{2}$ we can use it to completely determine $T$, and hence, an Einstein connection $\nabla$.
@@
Let $(M^{2n},f,g)$ be a weak almost Hermitian manifold, considered as a nonsymmetric pseudo-Riemannian manifold $(M,G=g+F)$, where $F(X,Y)=g(X,fY)$. Suppose that an Einstein connection $\nabla$ on $M$ satisfies the $f^{2}$-torsion condition (3.3), and the tensor $P:={I}-f^{2}$ has rank $2n$. Then the torsion $T$ of $\nabla$ is given by
$2\,T(Y,Z,\,(I+\tfrac{1}{2}\,\widetilde{Q})^{2}f^{4}X)=(\nabla^{g}_{X}F)((f+f^{3})Y,fZ)+(\nabla^{g}_{Y}F)(f^{2}Z,X)+(\nabla^{g}_{Z}F)(f^{2}X,Y)$
$\qquad+(\nabla^{g}_{fX}F)(f^{3}Y,Z)-(\nabla^{g}_{fX}F)(f^{2}Y,fZ)+(\nabla^{g}_{fY}F)(f^{3}Z,X)+(\nabla^{g}_{fZ}F)(f^{2}X,fY)$
$\qquad+(\nabla^{g}F)(P^{-1}(2\,\widetilde{Q}+\widetilde{Q}^{2})f^{2}X,fY,Z)-(\nabla^{g}F)(P^{-1}(2\,\widetilde{Q}+\widetilde{Q}^{2})f^{2}X,Y,fZ)$
$\qquad-(\nabla^{g}F)(\widetilde{Q}f^{2}X,Y,Z)-\tfrac{1}{2}\,dF((3\,\widetilde{Q}+2\,\widetilde{Q}^{2})X,fY,fZ)-\tfrac{3}{2}\,dF(\widetilde{Q}f^{4}X,Y,Z).$ (4.5)
@@
Applying $Y\to fY$ and $Z\to fZ$ to (2.6), we have